# Generative AI Agents for Satellite Networks through a Mixture of Experts Transmission

### v1

@@ I Introduction
  • Challenge II: Compared with ground communication systems, satellite communication networks also face more complex challenges in resource allocation [8]. Due to the wide coverage of satellite communications and the large space where users are distributed, coupled with the limited resources and fixed beam coverage design of satellites, it is difficult to meet the diverse and varying requirements of different users [9]. Moreover, transmission interference between satellites, between beams, and between users seriously affects communication performance [10]. Therefore, developing resource allocation schemes to mitigate interference and manage resources effectively is critical to improving the quality of service (QoS) of satellite communication networks.
@@
To address Challenge II, the mixture of experts (MoE) model has been introduced. MoE integrates multiple specialized deep neural network (DNN) components, i.e., experts, to handle complex tasks [16]. Specifically, each expert is fine-tuned and specialized for a specific type of sub-task or sub-dataset, providing expertise in efficient problem-solving capabilities within their domain [17]. MoE uses a gating network technique to divide larger tasks into smaller, more manageable units and dynamically select the most appropriate combination of experts to handle these tasks based on the input data characteristics. Unlike traditional neural network models, MoE selectively activates a subset of experts based on specific features of the input, thereby improving the model’s processing efficiency and performance [18]. In satellite communication networks, MoE can provide customized solutions for various challenges, such as spectrum allocation, signal interference reduction, and energy efficiency optimization. For example, under dynamically changing communication environments and diverse requirements of different users, MoE can achieve effective resource allocation and interference management by selecting a combination of experts that are most suitable for handling current signal conditions and user distribution [19].
Inspired by the above discussion, as shown in Fig. 1, this paper first studies a framework for modeling using generative AI agents and then optimizes the transmission strategy of satellite communication networks through an MoE model. To the best of the authors’ knowledge, this is the first work in the networking field to adopt generative AI agents for customized modeling and then solve the formulated problem by MoE. [^1] https://github.com/RickyZang/GAI-agent-satellite The contributions of this work are summarized as follows.
@@
  • Based on different satellite scenarios, we employ different access strategies, channel models, and optimization objectives for heterogeneous and homogeneous satellite networks through the generative AI agent framework to customize our system modeling. This step ensures that the models accurately reflect the complexity and specific requirements of the actual communication environment in both LEO and Geostationary Earth Orbit (GEO) satellite networks. (For Challenge I)
  • We propose an MoE-based proximal policy optimization (PPO) method to optimize the formulated problem generated by the generative AI agent. Specifically, each expert can optimize the optimization variables it is good at through its network and then joint aggregate them through the gating network to perform joint optimization, which can maximize the efficiency of spectrum resource utilization while ensuring communication quality. (For Challenge II)
@@ II-A Satellite Communication Networks
Recent advancements in satellite communication networks have garnered significant attention due to their potential benefits of global connectivity. For instance, in [20], an optimization problem for maximizing the sum rate in massive MIMO LEO satellite networks, considering both imperfect successive interference cancellation (SIC) and channel state information (CSI). Similarly, Huang et. al. [21] addressed the sum rate maximization problem in LEO satellite-terrestrial networks utilizing RSMA and proposed a PPO-based approach to meet diverse QoS requirements. Khan et. al. [22] focused on sum rate maximization in heterogeneous satellite networks, introducing efficient subcarrier beam assignment and power allocation strategies for secondary LEO satellites while managing interference with primary GEO satellites. Besides spectral efficiency, in [23], the system EE was maximized in integrated satellite-terrestrial radio access networks by jointly optimizing the cache sharing vector, block placement vector, and power allocation vector, where the information EH QoS requirements of users and the power budget are considered as constraints. Despite these contributions, the complex models inherent in these efforts may create barriers to understanding for newcomers to the field, which requires a framework that makes entry into this field of research more accessible.
@@ II-B Generative AI Agent
The field of generative AI has seen significant interest, particularly in its application to interactive agents. For instance, Wang et. al. [24] developed an interactive chatbot leveraging LLMs to mimic character-specific personality traits, achieving a high correlation (i.e., about 82.8%) with human perception. Also, The SWIFTSAGE framework, introduced by Lin et. al. [25], combines behavioral cloning with LLM prompting for enhanced action planning in complex tasks. Furthermore, generating agents are also used in the networking field. For instance, Du et. al. [13] demonstrated a generative AI agent empowered by LLMs to simulate user interactions, offering insights into real-time Quality of Experience (QoE) feedback, where experiments showed that using the proposed agent can improve performance by 15%. Also, Zhang et. al. [12] proposed an optimization framework including environment, action, brain, and perception by using an Generative AI Agent, where pluggable LLM and RAG modules are employed for decision-making knowledge base and contextual memory. However, these aforementioned works focus mainly on direct policy generation and do not consider the potential for customized and selective system modeling and network configuration.
@@ III-A3 Channel Models
Channel modeling in satellite communication models is quite different, such as fixed and time-varying channels [33]. For the fixed channel model, it assumes that the channel characteristics remain unchanged over a long time, where the corresponding channel model is expressed by
@@
For the time-varying channel model, it is typically that channel characteristics such as Doppler shift change rapidly with time. Following Jakes model [34], keeping the large-scale fading in (9) unchanged, the small-scale fading vector g is modeled as a first-order complex Gauss Markov process, i.e.,
@@ III-A4 Optimization Goals
To this end, a generative AI agent framework is constructed, as shown in Fig. 2. Specifically, first, we utilize LLMs to establish a conversational modeling procedure. In each round of interaction, the semantic router [36] integrated into LLM extracts task-relevant semantics from users’ natural language and invokes the corresponding function calls. Since LLMs are trained on general knowledge and lack satellite expertise, we build an RAG [37] system, which contains massive expert knowledge regarding satellite communications to support complicated mathematical modeling. The RAG functions called by the semantic router retrieve relevant knowledge from the dataset. Afterward, the agent leverages LLMs to analyze the retrieved knowledge and generate the configurations of each modeling aspect step-by-step, based on user descriptions and requirements. Next, we introduce the detailed design.
@@ III-B Semantic Router
As shown in Fig. 2, we segment the entire expertise dataset into four blocks, each of which corresponds to one aspect mentioned in Section III-A. Moreover, each block contains two sub-blocks, corresponding to the specific configurations in each aspect. In this case, to generate accurate system models, the prerequisite for our generative agent is to locate the correct sub-block for expertise retrieval. The semantic router acts as the nerves of agents, transforming users’ descriptions in natural language to the calling for the specific sub-block. To do so, it employs a two-layer structure, in which layer-1 routing realizes the mapping between user description and block. Afterward, the layer-2 routing further determines which sub-block to call. The two routing selection layers are based on semantic similarity [36]. Specifically, denoting user description as $t$, the semantic router will use an encoder $\mathcal{E}$ to encode it as text embeddings $\mathbf{e}(t)$. The labels of blocks and sub-blocks are $\mathbf{b}_{i}$ and ($\mathbf{b}_{i}$, $\mathbf{s}^{i}_{j}$), where $\mathbf{b}_{i}$ and $\mathbf{s}^{i}_{j}$ ($i\in\{1,2,3,4\},j\in\{1,2\}$) refer to the aspect and configuration, respectively. For example, $\mathbf{b}_{3}$ = “channel model" and $\mathbf{s}^{3}_{1}$ = “SDMA" (see Fig. 2). Using cosine similarity as the metric for similarity measurement, the layer-1 routing can be defined as follows
$\min_{i}\left(1.0-\frac{\sum\limits_{k=1}^{\zeta}\left(\mathbf{e}(t)[k]\times\mathbf{e}(\mathbf{b}_{i})[k]\right)}{\sqrt{\sum\limits_{k=1}^{\zeta}\left(\mathbf{e}(t)[k]\right)^{2}}\sqrt{\sum\limits_{k=1}^{\zeta}\left(\mathbf{e}(\mathbf{b}_{i})[k]\right)^{2}}}\right),$ (19)
where $\zeta$ denotes the length of semantic embeddings. [$k$] means the $k$-th element of any given embedding vector. Similarly, layer-2 routing follows the same principle, while it aims at finding $j\in\{1,2\}$ which minimizes the cosine distance between $\mathbf{e}(t)$ and $\mathbf{e}\left((\mathbf{b}_{i},\mathbf{s}^{i}_{j})\right)$.
@@ III-C Retrieval-Augmented Generation
After mapping user descriptions to the specific RAG sub-block, the next operation is to retrieve relevant expertise knowledge from the dataset to support satellite communications modeling. Such a process can be modeled as using input sequence $q$ to fetch knowledge $z$, subsequently utilized as supplementary context for helping LLMs to generate the target sequence $d$. Inspired by [37], we develop our RAG system atop RAG-Sequence, utilizing identical knowledge to forecast each token within the target sequence. As illustrated in Fig. 2, our system encompasses two pivotal elements: (i) a retrieval mechanism $p_{\eta}(z|q)$ characterized by parameters $\eta$, tasked with deriving a truncated distribution over knowledge embeddings in response to a query $q$, and (ii) a generative module $p_{\theta}(d_{i}|q,z,d_{1:i-1})$, parameterized by $\theta$, responsible for the sequential generation of tokens, each contingent upon the antecedent $i-1$ tokens $d_{1:i-1}$, the initial query $q$, and retrieved knowledge $z$. Such a process encompasses the marginalization of the seq2seq probability $p_{\rm Seq}(d|x)$ via a top-K approximation [37], formalized as follows:
$p_{\rm Seq}(d|q)$ $\approx\sum_{z\in{\rm top-k}(p(\cdot|q))}p_{\eta}(z|q)p(d|z,q)$ (20)
$=\sum_{z\in top-k(p(\cdot|q))}p_{\eta}(z|q)\prod_{i}^{L}p_{\theta}(d_{i}|q,z,d_{1:i-1}),$
@@ III-C1 Retrieval Mechanism
$p_{\eta}(z|q)\propto\exp\left(\mathbf{E}_{\rm t}(z)^{\top}\mathbf{E}_{\rm q}(q)\right),$ (21)
where $\mathbf{E}_{\rm t}(z)$ is a dense representation of the knowledge produced by the aforementioned text encoder [38], and $\mathbf{E}_{\rm q}(q)$ denotes the query representation produced by a query encoder. Due to its strong representation ability, such encoders are now usually implemented on top of LLM, such as text-embedding-ada-002 [39], i.e.,
$\mathbf{E}_{\rm t}(z)=\text{LLM}_{e}(z)\;\;\&\;\;\mathbf{E}_{\rm q}(q)=\text{LLM}_{q}(q).$ (22)
@@ III-C2 Generative Module
The generative module is pluggable and can be served by any mainstream LLM. Here, we adopt a generative pre-trained transformer (GPT) to realize $p_{\theta}(d_{i}|q,z,d_{1:i-1})$, organizing retrieved knowledge and generating coherent and context-aware answers about satellite communications modeling.
@@ III-C3 Training Approaches
The training scheme jointly optimizes the retriever and generator without explicit directives on document retrieval. The objective minimizes the negative marginal log-likelihood of the target sequences across a corpus of input/output pairs $(q_{j},d_{j})$, employing stochastic gradient descent Adam.

### v2

# Generative AI Agents with Large Language Model for Satellite Networks via a Mixture of Experts Transmission

@@ I Introduction
  • Challenge II:Compared with ground communication systems, satellite communication networks face more complex challenges in resource allocation [8]. Due to the wide coverage of satellite communications and the large space where users are distributed, coupled with the limited resources and fixed beam coverage design of satellites, it is difficult to meet the diverse and varying requirements of different kinds of users [9]. Moreover, transmission interference between satellites, between beams, and between users seriously affects communication performance [10]. Therefore, developing adaptive resource allocation schemes to mitigate interference and manage resources effectively is critical to improving the quality of service (QoS) of satellite communication networks.
@@
To address Challenge II, the mixture of experts (MoE) model has been introduced. MoE integrates multiple specialized deep neural network (DNN) components, i.e., experts, to handle complex tasks [16]. Specifically, each expert is fine-tuned and specialized for a specific type of sub-task or sub-dataset, providing expertise in efficient problem-solving capabilities within their domain, which can be more effective than monolithic models [17]. MoE uses a gating network technique to divide larger tasks into smaller, more manageable units and dynamically select the most appropriate combination of experts to handle these tasks based on the input data characteristics. Unlike traditional neural network models, MoE selectively activates a subset of experts based on specific features of the input, thereby improving the model’s processing efficiency and performance [18]. In satellite communication networks, MoE can provide customized solutions for various challenges based on different experts, such as spectrum allocation, signal interference reduction, and energy efficiency optimization. For example, under dynamically changing communication environments and diverse requirements of different users, MoE can achieve effective resource allocation and interference management by selecting a combination of experts that are most suitable for handling current signal conditions and user distribution [19].
Inspired by the above discussion, as shown in Fig. 1, this paper first studies a framework for modeling using generative AI agents and then optimizes the transmission strategy of satellite communication networks through an MoE model. For the generative AI agents, unlike traditional system modeling, which typically relies on the designer’s knowledge and experience to manually formulate mathematical models, our generative AI agents leverage the vast knowledge embedded in LLMs and the precise contextual information provided by RAG to customize and adapt the modeling process to specific satellite communication needs. For the transmission strategy, instead of using a single comprehensive solution system to optimize all variables or employing a multi-agent system to address different aspects, our proposed MoE approach utilizes multiple specialized experts, each focusing on optimizing specific variables, thereby improving efficiency. To the best of the authors’ knowledge, this is the first work in the networking field to adopt generative AI agents for customized modeling and then solve the formulated problem by MoE. [^1] https://github.com/RickyZang/GAI-agent-satellite The contributions of this work are summarized as follows.
@@
  • Based on different satellite scenarios, we employ different access strategies, channel models, and optimization objectives for heterogeneous and homogeneous satellite networks through the generative AI agent framework to customize our system modeling. This step ensures that the models accurately reflect the complexity and specific requirements of the actual communication environment in both LEO and Geostationary Earth Orbit (GEO) satellite networks. By doing so, we enable a more efficient selection of modeling parameters based on designers’ requirements, reducing the risk of errors and saving time. (For Challenge I)
  • We propose an MoE-based proximal policy optimization (PPO) method to optimize the formulated problem generated by the generative AI agent. Specifically, each expert can optimize the optimization variables for which it has been trained through its network and then aggregate them through the gating network to perform joint optimization, which can maximize the efficiency of spectrum resource utilization while meeting communication quality requirements. (For Challenge II)
@@ II-A Satellite Communication Networks
Recent advancements in satellite communication networks have garnered significant attention due to their potential benefits of global connectivity. For instance, Chen et. al. [3] proposed federated learning in LEO satellite communication networks to support massively interconnected devices and reduce communication overheads, showing through simulations that the method significantly reduces latency and improves satellite communication performance in 6G systems. Also, in [20], an optimization problem for maximizing the sum rate in massive MIMO LEO satellite networks, considering both imperfect successive interference cancellation (SIC) and channel state information (CSI). Similarly, Huang et. al. [21] addressed the sum rate maximization problem in LEO satellite-terrestrial networks utilizing RSMA and proposed a PPO-based approach to meet diverse QoS requirements. Khan et. al. [22] focused on sum rate maximization in heterogeneous satellite networks, introducing efficient subcarrier beam assignment and power allocation strategies for secondary LEO satellites while managing interference with primary GEO satellites. Besides spectral efficiency, in [23], the system EE was maximized in integrated satellite-terrestrial radio access networks by jointly optimizing the cache sharing vector, block placement vector, and power allocation vector, where the information EH QoS requirements of users and the power budget are considered as constraints. Despite these contributions, the complex models inherent in these efforts may create barriers to understanding for newcomers to the field, which requires a framework that makes entry into this field of research more accessible.
@@ II-B Generative AI Agent
The field of generative AI has seen significant interest, particularly in its application to interactive agents [24]. For instance, Wang et. al. [25] developed an interactive chatbot leveraging LLMs to mimic character-specific personality traits, achieving a high correlation (i.e., about 82.8%) with human perception. Also, The SWIFTSAGE framework, introduced by Lin et. al. [26], combines behavioral cloning with LLM prompting for enhanced action planning in complex tasks. Furthermore, generating agents are also used in the networking field. For instance, Du et. al. [13] demonstrated a generative AI agent empowered by LLMs to simulate user interactions, offering insights into real-time Quality of Experience (QoE) feedback, where experiments showed that using the proposed agent can improve performance by 15%. Also, Zhang et. al. [12] proposed an optimization framework including environment, action, brain, and perception by using an Generative AI Agent, where pluggable LLM and RAG modules are employed for decision-making knowledge base and contextual memory. However, these aforementioned works focus mainly on direct policy generation and do not consider the potential for customized and selective system modeling and network configuration.
@@ III-A3 Channel Models
Channel modeling in satellite communication models is quite different, such as fixed and time-varying channels [34]. For the fixed channel model, it assumes that the channel statistical characteristics remain unchanged over a long time, where the corresponding channel model is expressed by
@@
For the time-varying channel model, it is typically that channel statistical characteristics such as Doppler shift change rapidly with time. Following Jakes model [35], keeping the large-scale fading in (9) unchanged, the small-scale fading vector g is modeled as a first-order complex Gauss Markov process, i.e.,
@@ III-A4 Optimization Goals
To this end, a generative AI agent framework is constructed, as shown in Fig. 2. Specifically, we first utilize LLMs to establish a conversational modeling procedure. In each round of interaction, the semantic router [37] integrated into LLM extracts task-relevant semantics from users’ natural language and invokes the corresponding function calls. Since LLMs are trained on general knowledge and lack satellite expertise, we build a RAG [14] system, which contains massive expert knowledge regarding satellite communications to support complicated mathematical modeling. The RAG functions called by the semantic router retrieve relevant knowledge from the dataset. Afterward, the agent leverages LLMs to analyze the retrieved knowledge and generate the configurations of each modeling aspect step-by-step, based on user descriptions and requirements. Next, we introduce the detailed design.
@@ III-B Semantic Router
As shown in Fig. 2, we segment the entire expertise dataset into four blocks, each corresponding to one aspect mentioned in Section III-A. Moreover, each block contains two sub-blocks, corresponding to the specific configurations in each aspect. In this case, to generate accurate system models, the prerequisite for our generative agent is to locate the correct sub-block for expertise retrieval. The semantic router acts as the nerves of agents, transforming users’ descriptions in natural language to the calling for the specific sub-block. To do so, it employs a two-layer structure, in which layer-1 routing realizes the mapping between user description and block. Afterward, the layer-2 routing further determines which sub-block to call. The two routing selection layers are based on semantic similarity [37]. Specifically, denoting a user description/query as $q\in\mathbf{q}$, the semantic router will use an encoder $\mathbf{E}_{q}$ to encode it as text embeddings, i.e., $\mathbf{E}_{q}(q)$. Here, text embeddings play a crucial role in bridging humans and LLMs by encoding natural language into a numerical format that neural networks can process. In this paper, we leverage the pre-trained text-embedding-ada-002 [38] model to generate embeddings of textual inputs. By training on large corpora, this model acquires outstanding contextual understanding capabilities, projecting each word/phrase into the continuous feature space with minimized conflicts. Moreover, we adopt the widely used cosine distance [37] to measure the semantic similarity between two embeddings.
The labels of blocks and sub-blocks are $\mathbf{b}_{i}$ and ($\mathbf{b}_{i}$, $\mathbf{s}^{i}_{j}$), where $\mathbf{b}_{i}$ and $\mathbf{s}^{i}_{j}$ ($i\in\{1,2,3,4\},j\in\{1,2\}$) refer to the aspect and configuration, respectively. For instance, $\mathbf{b}_{3}$ = “channel model”, $\mathbf{s}^{3}_{1}$ = “SDMA”, and $\mathbf{s}^{3}_{2}$ = “RSMA” (see Fig. 2). Using cosine similarity as the metric for similarity measurement, the layer-1 routing of query $q$ can be defined as follows:
$\min_{i}\left(1.0-\frac{\sum\limits_{k=1}^{\zeta}\left(\mathbf{E}_{q}(q)[k]\times\mathbf{E}_{q}(\mathbf{b}_{i})[k]\right)}{\sqrt{\sum\limits_{k=1}^{\zeta}\left(\mathbf{E}_{q}(q)[k]\right)^{2}}\sqrt{\sum\limits_{k=1}^{\zeta}\left(\mathbf{E}_{q}(\mathbf{b}_{i})[k]\right)^{2}}}\right),$ (19)
where $\zeta$ denotes the length of text embeddings. From this equation, we can observe that the user query will be routed to the block whose label is the most relevant. Similarly, layer-2 routing repeats Eq. (19) while aiming to further route $q$ to the most relevant sub-block.
Remark 2.
The semantic router is presented to enhance the efficiency of RAG for satellite communications modeling. This is because the rate of correct retrieval decreases with the increasing knowledge volume. By organizing expertise knowledge into eight sub-blocks, the space for retrieval can be reduced by eight times, making our proposal suitable for complicated satellite communication modeling. Moreover, we note that the block and sub-block configuration is customizable. Users can include/remove blocks/sub-blocks according to the specific application scenarios. After routing the user query to the specific sub-block, the RAG-assisted modeling can be performed, which is discussed in the next part.
@@ III-C Retrieval-Augmented Generation
RAG aims to retrieve relevant expertise from the knowledge sub-block to support satellite communications modeling. Such a process can be modeled as using input sequence $q$ to fetch knowledge $z$, subsequently utilized as supplementary context for helping LLMs to generate the modeling answers $m$. Inspired by [14], we develop our RAG system atop RAG-Sequence, utilizing identical knowledge to forecast each token within the target sequence. As illustrated in Fig. 2, our system encompasses two pivotal elements: (i) a retrieval mechanism $p_{\eta}(z|q)$ characterized by parameters $\eta$, tasked with deriving a truncated distribution over knowledge embeddings in response to a query $q$, and (ii) a generative module $p_{\theta}(m_{i}|q,z,m_{1:i-1})$, parameterized by $\theta$, responsible for the sequential generation of tokens, each contingent upon the antecedent $i-1$ tokens $m_{1:i-1}$, the initial query $q$, and retrieved knowledge $z$. Such a process encompasses the marginalization of the seq2seq probability $p_{\rm Seq}(m|x)$ via a top-K approximation [14], formalized as follows:
$p_{\rm Seq}(m|q)$ $\approx\sum_{z\in{\rm top-K}(p(\cdot|q))}p_{\eta}(z|q)p(m|z,q)$ (20)
$=\sum_{z\in top-K(p(\cdot|q))}p_{\eta}(z|q)\prod_{i}^{L}p_{\theta}(m_{i}|q,z,m_{1:i-1}),$
@@
1 $\mathbf{Input:}$ User query, embedding models, LLM;
2 Initialize environment;
3 Listen to user query $\mathbf{q}=\{q_{1},q_{2},\dots,q_{n}\}$
4 for each $q$ in $\mathbf{q}$ do
5       Encode $q$ and acquire embeddings $\mathbf{E}_{\rm q}(q)$;
6       for each $i$ in {1, 2, 3, 4} do
7             Calculate cosine similarity between $\mathbf{E}_{\rm q}(q)$ and $\mathbf{E}_{\rm q}(\mathbf{b}_{j})$ based on (19);
8
9      Route to the corresponding block $i$; # layer-1 routing
10       for each $j$ in {1, 2} do
11             Calculate cosine similarity between $\mathbf{E}_{\rm q}(q)$ and $\mathbf{E}_{\rm q}(\mathbf{s}^{i}_{j})$ based on (19);
12
13      Route to the corresponding sub-block $j$; # layer-2 routing
14       for each chunk in the sub-block do
15             Calculate cosine similarity between $\mathbf{E}_{\rm q}(q)$ and the chunk based on (19);
16
17      Fetch the most prevalent knowledge chunks;
18       Call LLM to generate answers for $q$ based on retrieved knowledge;
19
$\mathbf{Output:}$ The entire satellite communications model containing scenario, access protocol, channel, and optimization goal;
Algorithm 1 Generative AI agent for satellite communications modeling
@@ III-C1 Retrieval Mechanism
$p_{\eta}(z|q)\propto\exp\left(\mathbf{E}_{\rm z}(z)^{\top}\mathbf{E}_{\rm q}(q)\right),$ (21)
where $\mathbf{E}_{\rm z}(z)$ is a dense representation of the knowledge produced by the aforementioned text encoder [39], and $\mathbf{E}_{\rm q}(q)$ denotes the query representation produced by a query encoder. Due to its strong representation ability, such encoders are now usually implemented on top of LLM, such as text-embedding-ada-002 [38], i.e.,
$\mathbf{E}_{\rm z}(z)=\text{LLM}_{e}(z)\;\;\&\;\;\mathbf{E}_{\rm q}(q)=\text{LLM}_{q}(q).$ (22)
The knowledge in the selected sub-block is split into multiple chunks. Based on Eq. (21), the most relevant chunks can be retrieved to support the generation of satellite communications modeling.
@@ III-C2 Generative Module
The generative module is pluggable and can be served by any mainstream LLM. Here, we adopt a generative pre-trained transformer (GPT) to realize $p_{\theta}(m_{i}|q,z,m_{1:i-1})$, organizing retrieved knowledge and generating coherent and context-aware answers about satellite communications modeling.
@@ III-C3 Training Approaches
The training scheme jointly optimizes the retriever and generator without explicit directives on document retrieval. The objective minimizes the negative marginal log-likelihood of the target sequences across a corpus of input/output pairs $(q_{j},m_{j})$, employing stochastic gradient descent Adam.